# Quadratic Modelings of Syndrome Decoding

### v1

@@ 1 Introduction
The parameters of many code-based schemes are carefully chosen to align with the latest advancements with respect to this computational problem. Despite decades of intensive research in this direction, all the algorithms developed so far exhibit exponential complexity. This is not surprising, since the problem has been shown to be NP-hard [8]. In particular, after more than 60 years of investigation since the groundbreaking paper of Prange [34], the reduction in the exponent for most parameters of interest has been minimal [37, 23, 27, 10, 29, 7, 30, 12]. All the works mentioned fall into the family of Information Set Decoding (ISD) algorithms, whose basic observation is that it is easier to guess error-free positions, and guessing enough of them is sufficient to decode. This resilience against ISD algorithms adds to the robustness of the syndrome decoding problem as the underlying assumption for code-based cryptosystems.
To comprehensively assess security, it is imperative to consider attacks stemming from various other realms of post-quantum cryptography. For instance, attacks typically associated with multivariate or lattice-based schemes should also be taken into account for code-based schemes, when applicable. A remarkable example is offered by dual attacks, originally introduced in lattice-based cryptography, but strongly questioned in that context. In contrast, their code-based counterpart [17, 18] has recently outperformed ISD techniques for a non-negligible regime of parameters, by reducing the decoding problem to the closely related Learning Parity with Noise problem. Concerning polynomial system solving strategies, another notable illustration of this is the algebraic MinRank attack, which broke the rank-metric code-based schemes RQC and Rollo [4, 5] and now represents the state-of-the-art for MinRank cryptanalysis, beating combinatorial approaches.
In the Hamming metric, a reduction that transforms an instance of the syndrome decoding problem into a system of quadratic equations over $\mathbb{F}_{q}$ was introduced in [33]. The goal of [33] was to provide explicit reductions between the Syndrome Decoding Problem (SDP) and the Multivariate Quadratic (MQ) problem, thus proving an isomorphism between these two NP-complete problems. Focusing on the reduction from MLD to MQ, the most expensive step of the transformation, in terms of numbers of new variables and new equations introduced, is the so-called Hamming-weight computation encoding. Indeed, for a binary linear code of length $n$, the procedure dominates the overall complexity of the reduction with a complexity of $\mathcal{O}(n\log_{2}(n)^{2})$.
@@
We also provide theoretical and experimental data to analyze the complexity of solving syndrome decoding instances using our modelings, demonstrating that, at least for small parameters, our new strategy is practical and successful. Software (MAGMA scripts) supporting the work in this chapter can be found here.
@@ 2 Preliminaries
This paper investigates the reduction of the Syndrome Decoding Problem into a Polynomial System Solving Problem (PoSSo). In this section, we briefly recall the definitions of both problems, as well as the notions of solving degree and degree of regularity, which are commonly used to estimate the computational complexity of the PoSSo problem.
@@ 2.1 The Syndrome Decoding Problem
An $[n,k]$-linear code $\mathcal{C}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$. We call $n$ the length of the code, and $k$ its dimension. An element $\mathbf{x}\in\mathbb{F}_{q}^{n}$ is called a codeword if $\mathbf{x}\in\mathcal{C}$. The number of nonzero elements in $\mathbf{x}$ is called the Hamming weight of $\mathbf{x}$ and we denote it as $\mathsf{wt}(\mathbf{x})$. Given a code $\mathcal{C}$ we define a parity check matrix of $\mathcal{C}$ as $\mathbf{H}\in\mathbb{F}_{q}^{(n-k)\times n}$ such that the right kernel of $\mathbf{H}$ is the code $\mathcal{C}$. The subspace spanned by the rows of $\mathbf{H}$ is called the dual code of $\mathcal{C}$. Many code-based cryptosystems rely on the hardness of solving the Syndrome Decoding Problem (SDP), see Problems 1 and 2 described below.
@@
In later sections, we will also refer to Problem 1 as the “Bounded Syndrome Decoding” Problem. In this paper, we will also consider the following variant of SDP.
@@
It can be shown that, as long as the number of errors is below half the Gilbert-Varshamov distance, the Syndrome Decoding problem typically has a unique solution. Moreover, the instances where the number of errors attains the GV distance are those supposed to be the most difficult.
@@ 2.2 The Polynomial System Solving Problem
Under suitable assumptions, the degree of regularity provides an upper bound for the solving degree [16, 35, 36]. Moreover, it is often assumed that the two values of the degree of regularity and the solving degree are close. Although this occurs in many relevant situations, there are examples where these two invariants can be arbitrarily far apart (see [15, 22, 11]). We will see in Section 5 that the degree of regularity of the system presented in Section 4.2 seems to yield a much higher value than the solving degree achieved during the Gröbner basis algorithm.
@@ 3 The MPS Modeling
The idea is to construct an algebraic system having a variety containing elements $(\mathbf{x}\mid\mathbf{y}_{1}\mid\cdots\mid\mathbf{y}_{n})\in\mathbb{F}_{2}^{n(\ell+1)}$ whose first $n$ coordinates represent an element $\mathbf{x}$ of $\mathbb{F}_{2}^{n}$ such that $\mathbf{H}\mathbf{x}^{\top}=\mathbf{s}^{\top}$. The remaining $n\ell$ coordinates are considered to be the concatenation of $n$ elements $\mathbf{y}_{i}\in\mathbb{F}_{2}^{\ell}$ where the elements of $\mathbf{y}_{i}$ represent the binary expansion of $\mathsf{wt}(\pi_{i}(\mathbf{x}))$ for every $i=1,\ldots,n$, with $\pi_{i}(\mathbf{x})=(x_{1},\dots,x_{i})$. By this definition, the list $\mathbf{y}_{n}$ represents the binary expansion of $\mathsf{wt}(\mathbf{x})$. The system finally enforces that $\mathbf{y}_{n}$ represents the binary expansion of an integer $t^{\prime}$ such that $t^{\prime}\leq t$. The elements of the variety corresponding to the ideal generated by this algebraic modeling are finally projected onto their first $n$ coordinates, revealing the solutions to the original syndrome decoding problem.
@@
  • Hamming weight computation encoding. This part of the modeling provides a set of polynomials that describe the binary encoding of $\mathsf{wt}(\pi_{i}(\mathbf{x}))$ for every $i=1,\ldots,n$ described above. The set of polynomials achieving this goal, is given by the union of the three following sets consisting of the $\ell+n-1$ polynomials in the sets
@@
    $f_{j}=(y_{n,j}+v_{j})\prod_{h=j+1}^{\ell}y_{n,h}+v_{h}+1$
    for $j=1,\ldots,\ell$. The output set is the singleton
@@ 4.1 Improved Modeling for the Case of SDP
We consider the $\mathsf{degrevlex}$ monomial ordering on $\mathbb{F}_{2}[X,Y]$ and denote by $\mathsf{lm}(p)$ the leading monomial of a polynomial $p$. Notice that since we are in the binary case, the notions of leading monomial and that of leading term coincide.
@@
Set now $\ell_{t}=\lfloor\log_{2}(t)\rfloor+1$. A further improvement to the MPS modeling (described in Equation 7) follows by observing that in the non-trivial case where $t<n$, we can impose that the last $\ell-\ell_{t}$ entries of $\mathbf{y}_{i}$ must be 00 for every $i=1,\ldots,n$. This means that we can add the linear equations $y_{i,j}=0$ for every $i=0,\ldots,n-1$ and $j=\ell_{t}+1,\ldots,\ell$. By inspection, setting the aforementioned variables to 00 will make part of the equations of the Hamming weight computation encoding vanish. We can equivalently simply consider the equations that remain, and get rid of the variables which have been set to 00. Consider the following updated notation.
@@
Let $n\geq 2$ and let $\mathcal{C}\subseteq\mathbb{F}_{2}^{n}$ be a $[n,k,d]$-linear code having a parity check matrix $\mathbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$. We define $\ell_{t}=\lfloor\log_{2}(t)\rfloor+1$. Let $\mathbf{s}\in\mathbb{F}_{2}^{n-k}$ play the role of the syndrome and let $0\leq t\leq\lfloor(d-1)/2\rfloor$ be the target error weight. Let $X=\left(x_{1},\ldots,x_{n}\right)$ and $Y=(Y_{1},\dots,Y_{n})$ with $Y_{j}=(y_{j,1},\dots,y_{j,\ell_{t}})$ be two sets of variables and consider the polynomial ring $\mathbb{F}_{2}[X,Y]$.
@@
The effect on the weight constraint encoding is simply the decrease in the degree from $\ell$ to $\ell_{t}$ of the produced polynomial. This is the only non-quadratic polynomial left in the modeling. We can turn this polynomial into a set of $\mathcal{O}(t\ell_{t})$ polynomials of degree up to $2$ in $\mathcal{O}(t\ell_{t})$ variables with the same linearization techniques described in [33, Fact 1 and Lemma 11].
@@
The improved modeling is an algebraic system of $\mathcal{O}(n(\ell_{t}+2)-k+t\ell_{t})$ polynomials of degree at most $2$ in $\mathcal{O}(n(\ell_{t}+1)+t\ell_{t})$ variables. Note that usual applications of the SDP to code-based cryptography choose $t\ll n$, hence the asymptotic bounds on the number of polynomials and variables in the improved modeling are both $\mathcal{O}(n\ell_{t})$. As shown in Table 1, our modeling improves over MPS by a factor of $\log_{2}(n)\log_{t}(n)$.
@@
Table 1: Comparison with the asymptotic size of the polynomial system in [33, Theorem 13], where $n$ is the length of the code and $t$ the bound on the weight of the target vector, that is $\mathsf{wt}(\mathbf{e})\leq t$.
@@ 5 Complexity Analysis of Modeling 2
We point out that the study in [13] investigates a particular case of the problem that this paper deals with, that is the regular syndrome decoding problem. The regular syndrome decoding problem considers error vectors having a regular distribution of non-zero entries. The algebraic modeling proposed in [13] has been conjectured to behave semi-regularly when the linear parity-check constraints and the fixed and structured quadratic polynomials are considered separately. Despite the fact that the problem tackled in [13] is a particular case of the problem we consider, our modeling has not been devised as a generalization of their modeling. Furthermore, we show that for the more general case, our modeling yields different results.
For the rest of this section, we retain the notation defined in Notation 2. Let $S\subset\mathbb{F}_{2}[X,Y]$ be the set of polynomials of Modeling 2 as described in Section 4.2. Let $L$ and $Q$ denote the sets of linear and quadratic polynomials, respectively. Clearly $S=L\cup Q$. Write also $L=L_{\mathbf{H}}\cup P$, where $L_{\mathbf{H}}$ denotes the set of linear polynomials in (2) introduced with the parity check matrix $\mathbf{H}$, and $P$ denotes the remaining linear polynomials in $S$. In other words, $P$ is the following set
$P=\left\{x_{1}+y_{1,1},y_{1,2},\ldots,y_{1,\ell_{t}}\right\}\cup\left\{x_{i}+y_{i,1}+y_{i-1,1}\mid i=2,\ldots,n\right\}.$
We want to estimate the degree of regularity of $S$. As mentioned in Section 2, this can be done by computing the degree of the Hilbert series of $\mathbb{F}_{2}[X,Y]/\langle S^{\mathrm{top}}\rangle$. Since we do not know $L_{\mathbf{H}}$ a priori, we first consider the set $S\setminus L_{\mathbf{H}}=Q\cup P$ and compute its degree of regularity. After, we will shortly argue the effect of adding $L_{\mathbf{H}}$ back to the set on the degree of regularity. We break down the problem even further by first computing the degree of regularity of $Q$ and then that of $Q\cup P$. We take advantage of the fact that the Hilbert series of $Q$ and of $Q\cup P$ are polynomials and compute their degree, i.e. for instance, $d_{\mathrm{reg}}(Q)=\deg\mathrm{HS}_{\mathbb{F}_{2}[X,Y]/\langle Q^{\mathrm{top}}\rangle}(z)+1$ as per Remark 2, similarly for $Q\cup P$. To this end, we are going to compute the maximum degree of a monomial in $\mathbb{F}_{2}[X,Y]/\langle Q^{\mathrm{top}}\rangle$, similarly we do for $Q\cup P$.
@@
and the ideal $\langle Q^{\mathrm{top}}\rangle\subseteq\mathbb{F}_{2}[X,Y]$ is thus a monomial ideal.
@@
  • If $n$ is even, there are exactly two maximal sets $\mathcal{S}\subseteq\mathcal{N}$ such that no set in $\mathcal{P}$ is a subset of $\mathcal{S}$.
  • If $n$ is odd, there is exactly one maximal set $\mathcal{S}\subseteq\mathcal{N}$ such that no set in $\mathcal{P}$ is a subset of $\mathcal{S}$.
@@
We aim to find the number of maximal sets $\mathcal{S}\subseteq\mathcal{N}$ such that no pair from $\mathcal{P}$ (i.e., no two consecutive elements) appears in $\mathcal{S}$. In order to avoid pairs of consecutive elements, we can only select non-consecutive elements from $\mathcal{N}$. To maximize the size of $\mathcal{S}$, we select every other element from $\mathcal{N}$. The size of such a maximal set $\mathcal{S}$ is:
$|\mathcal{S}|=\left\lceil\frac{n}{2}\right\rceil.$
Thus:
  • If $n$ is even, a maximal set contains $\frac{n}{2}$ elements.
  • If $n$ is odd, the maximal set contains $\frac{n+1}{2}$ elements.
@@
    Start with 1 and select every other element:
    $\mathcal{S}_{1}=\{1,3,5,\dots,n-1\}.$
    This set contains all the odd-numbered elements of $\mathcal{N}$, and its size is $k$.
@@
    Start with 2 and select every other element:
    $\mathcal{S}_{2}=\{2,4,6,\dots,n\}.$
    This set contains all the even-numbered elements of $\mathcal{N}$, and its size is also $k$.
Since there are no other ways to select $k$ elements without picking consecutive elements, these are the only two maximal sets for $n$ even.
Case 2: $n$ is odd. Let $n=2k+1$. The largest possible set $\mathcal{S}$ will contain $k+1=\frac{n+1}{2}$ elements. In this case, there is only one way to construct a set of size $k+1$ that avoids consecutive elements, i.e. start with 1 and select every other element:
$\mathcal{S}_{1}=\{1,3,5,\dots,n\}.$
This set contains $k+1$ elements and avoids consecutive pairs. If we were to start with 2 and select every other element, we would only get $k$ elements:
$\mathcal{S}_{2}=\{2,4,6,\dots,n-1\}.$
This is not maximal, as it contains fewer than $k+1$ elements. Thus, for $n$ odd, there is exactly one maximal set. ∎
Lemma 2 can be used to prove the following corollary, which we will use to construct a maximal degree monomial in $\mathbb{F}_{2}[X,Y]/\langle Q^{\mathrm{top}}\rangle$.
@@
Let $n=8$ and $\ell_{t}=3$. According to Theorem 5.2 the degree of regularity of $Q$ is
$d_{\mathrm{reg}}(Q)=8+3\left\lceil\frac{8}{2}\right\rceil+1=21.$
Using MAGMA, we compute and report the Hilbert series of the quotient ring $\mathbb{F}_{2}[X,Y]/\langle Q^{\mathrm{top}}\rangle$, i.e.
@@
$\tilde{Q}^{\mathrm{top}}_{j,n}:=Q^{\mathrm{top}}_{j,n-1}\setminus\{y_{1,j}y_{2,j}\}\subset\mathbb{F}_{2}[y_{2,j},\ldots,y_{n-1,j}].$
Let $G$ be a Gröbner basis of $\langle Q^{\mathrm{top}}\cup P^{\mathrm{top}}\rangle$ as in Lemma 3. Due to the presence of the linear monomials contributed by $P^{\mathrm{top}}$ we observe that
@@
Let $n=8$ and $\ell_{t}=3$. According to Theorem 5.2 the degree of regularity of $Q\cup P$ is
$d_{\mathrm{reg}}(Q\cup P)=\left\lceil\frac{7}{2}\right\rceil+(3-1)\left\lceil\frac{6}{2}\right\rceil+1=11.$
Using MAGMA, we compute and report the Hilbert series of the quotient ring $\mathbb{F}_{2}[X,Y]/\langle Q^{\mathrm{top}}\cup P^{\mathrm{top}}\rangle$, i.e.

### v2

@@ 1 Introduction
The parameters of many code-based schemes are carefully chosen to align with the latest advancements with respect to this computational problem. Despite decades of intensive research in this direction, all the algorithms developed so far exhibit exponential complexity. This is not surprising, since the problem has been shown to be NP-hard [8]. In particular, after more than 60 years of investigation since the groundbreaking paper of Prange [35], the reduction in the exponent for most parameters of interest has been minimal [38, 23, 28, 10, 30, 7, 31, 12]. All the works mentioned fall into the family of Information Set Decoding (ISD) algorithms, whose basic observation is that it is easier to guess error-free positions, and guessing enough of them is sufficient to decode. This resistance to ISD algorithms makes the syndrome decoding problem a reliable foundation for code-based cryptosystems.
To comprehensively assess security, it is imperative to consider attacks stemming from various other realms of post-quantum cryptography. For instance, attacks typically associated with multivariate or lattice-based schemes should also be taken into account for code-based schemes, when applicable. A remarkable example is offered by dual attacks, originally introduced in lattice-based cryptography, where, however, they have been strongly questioned. In contrast, their code-based counterpart [17, 18] has recently outperformed ISD techniques for a non-negligible regime of parameters, by reducing the decoding problem to the closely related Learning Parity with Noise problem. Concerning polynomial system solving strategies, another notable illustration of this is the algebraic MinRank attack, which broke the rank-metric code-based schemes RQC and Rollo [4, 5] and now represents the state-of-the-art for MinRank cryptanalysis, beating combinatorial approaches.
In the Hamming metric, a reduction that transforms an instance of the syndrome decoding problem into a system of quadratic equations over $\mathbb{F}_{2}$ was introduced in [34]. The most expensive step of the transformation, in terms of numbers of new variables and new equations introduced, is the so-called Hamming-weight computation encoding. Indeed, for a binary linear code of length $n$, the procedure dominates the overall complexity of the reduction with a complexity of $\mathcal{O}(n\log_{2}(n)^{2})$.
@@
We also provide theoretical and experimental data to analyze the complexity of solving syndrome decoding instances using our modelings, demonstrating that, at least for small parameters, our new strategy is practical and successful. Software (MAGMA scripts) supporting this work can be found here.
@@ 2 Preliminaries
This paper investigates the reduction of the Syndrome Decoding Problem (SDP) into a Polynomial System Solving Problem (PoSSo). In this section, we briefly recall the definitions of both problems, as well as the notions of solving degree and degree of regularity, which are commonly used to estimate the computational complexity of the PoSSo problem.
@@ 2.1 The Syndrome Decoding Problem
An $[n,k]$-linear code $\mathcal{C}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$. We call $n$ the length of the code, and $k$ its dimension. An element $\mathbf{x}\in\mathbb{F}_{q}^{n}$ is called a codeword if $\mathbf{x}\in\mathcal{C}$. The number of nonzero entries in $\mathbf{x}$ is called the Hamming weight of $\mathbf{x}$ and we denote it as $\mathsf{wt}(\mathbf{x})$. Given a code $\mathcal{C}$ we define a parity check matrix of $\mathcal{C}$ as $\mathbf{H}\in\mathbb{F}_{q}^{(n-k)\times n}$ such that the right kernel of $\mathbf{H}$ is the code $\mathcal{C}$. The subspace spanned by the rows of $\mathbf{H}$ is called the dual code of $\mathcal{C}$. Many code-based cryptosystems rely on the hardness of solving the Syndrome Decoding Problem (SDP), see Problems 1 and 2 described below.
@@
In later sections, we will also refer to Problem 1 as the “Bounded Syndrome Decoding” Problem. We will also consider the following variant of SDP.
@@
It can be shown that, as long as the number of errors is below the Gilbert-Varshamov distance, the Syndrome Decoding problem typically has a unique solution. Moreover, the instances where the number of errors attains the GV distance are those supposed to be the most difficult.
@@ 2.2 The Polynomial System Solving Problem
Remark 2
If $\mathcal{F}$ is not homogeneous, Gaussian elimination on $M_{\leq d}$ may produce a row corresponding to a polynomial $f$ with $\deg f<d$, where the leading term of $f$ was not the leading term of any row in $M_{\leq d}$. Some algorithms, for example $F4$, address this by adding rows for polynomials $mf$ ($\deg(mf)\leq d$) for some monomial $m$ and recomputing the reduced row echelon form. If no Gröbner basis is found in degree $\leq d$, they proceed to higher degrees, potentially enlarging the span of $M_{\leq d}$ and reducing the solving degree. Throughout this paper, we consider only the case where no extra rows are added. Note that the solving degree as defined above is an upper bound on the degree at which algorithms using this variation terminate.
@@
Under suitable assumptions, the degree of regularity provides an upper bound for the solving degree [16, 36, 37]. Moreover, it is often assumed that the two values are close. Although this occurs in many relevant situations, there are examples where these two invariants can be arbitrarily far apart (see [15, 22, 11]). We will see in Section 5 that the degree of regularity of the system presented in Section 4.2 seems to yield a much higher value than the solving degree achieved during the Gröbner basis algorithm.
@@ 3 The MPS Modeling
The idea is to construct an algebraic system having a variety containing elements $(\mathbf{x}\mid\mathbf{y}_{1}\mid\cdots\mid\mathbf{y}_{n})\in\mathbb{F}_{2}^{n(\ell+1)}$ whose first $n$ entries represent an element $\mathbf{x}$ of $\mathbb{F}_{2}^{n}$ such that $\mathbf{H}\mathbf{x}^{\top}=\mathbf{s}^{\top}$. The remaining $n\ell$ entries are considered to be the concatenation of $n$ elements $\mathbf{y}_{i}\in\mathbb{F}_{2}^{\ell}$ where the elements of $\mathbf{y}_{i}$ represent the binary expansion of $\mathsf{wt}(\pi_{i}(\mathbf{x}))$ for every $i=1,\ldots,n$, with $\pi_{i}(\mathbf{x})=(x_{1},\dots,x_{i})$. By this definition, the list $\mathbf{y}_{n}$ represents the binary expansion of $\mathsf{wt}(\mathbf{x})$. The system finally enforces that $\mathbf{y}_{n}$ represents the binary expansion of an integer $t^{\prime}$ such that $t^{\prime}\leq t$. The elements of the variety of solutions of this algebraic modeling are finally projected onto their first $n$ coordinates, revealing the solutions to the original syndrome decoding problem.
@@
  • Hamming weight computation encoding. This part of the modeling provides a set of polynomials that describes the binary encoding of $\mathsf{wt}(\pi_{i}(\mathbf{x}))$ for every $i=1,\ldots,n$ described above. The set of polynomials achieving this goal, is given by the union of the three following sets consisting of the $\ell+n-1$ polynomials in the sets
@@
    $f_{j}=(y_{n,j}+v_{j})\prod_{h=j+1}^{\ell}(y_{n,h}+v_{h}+1)$
    for $j=1,\ldots,\ell$. The set is the singleton
@@ 4.1 Improved Modeling for the Case of SDP
We consider the $\mathsf{degrevlex}$ monomial ordering on $\mathbb{F}_{2}[X,Y]$ with the $X$ variables greater than the $Y$ variables, and denote by $\mathsf{lm}(p)$ the leading monomial of a polynomial $p$. Notice that since we are in the binary case, the notions of leading monomial and that of leading term coincide.
@@
Set now $\ell_{t}=\lfloor\log_{2}(t)\rfloor+1$. A further improvement to the MPS modeling (described in Equation (7)) follows by observing that in the non-trivial case where $t<n$, we can impose that the last $\ell-\ell_{t}$ entries of $\mathbf{y}_{i}$ must be 00 for every $i=1,\ldots,n$. This means that we can add the linear equations $y_{i,j}=0$ for every $i=1,\ldots,n$ and $j=\ell_{t}+1,\ldots,\ell$. By inspection, setting the aforementioned variables to 00 will make part of the equations of the Hamming weight computation encoding vanish. We can equivalently simply consider the equations that remain, and get rid of the variables which have been set to 00. Consider the following updated notation.
@@
Let $n\geq 2$ and let $\mathcal{C}\subseteq\mathbb{F}_{2}^{n}$ be a $[n,k,d]$-linear code having a parity check matrix $\mathbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$. Let $\mathbf{s}\in\mathbb{F}_{2}^{n-k}$ play the role of the syndrome and let $0\leq t\leq\lfloor(d-1)/2\rfloor$ be the target error weight. We define $\ell_{t}=\lfloor\log_{2}(t)\rfloor+1$. Let $X=\left(x_{1},\ldots,x_{n}\right)$ and $Y=(Y_{1},\dots,Y_{n})$ with $Y_{j}=(y_{j,1},\dots,y_{j,\ell_{t}})$ be two sets of variables and consider the polynomial ring $\mathbb{F}_{2}[X,Y]$.
@@
The effect on the weight constraint encoding is simply the decrease in the degree from $\ell$ to $\ell_{t}$ of the produced polynomial. This is the only non-quadratic polynomial left in the modeling. We can turn this polynomial into a set of $\mathcal{O}(t\ell_{t})$ polynomials of degree up to $2$ in $\mathcal{O}(t\ell_{t})$ variables with the same linearization techniques described in [34, Fact 1 and Lemma 11].
@@
The improved modeling is an algebraic system of $\mathcal{O}(n(\ell_{t}+2)-k+t\ell_{t})$ polynomials of degree at most $2$ in $\mathcal{O}(n(\ell_{t}+1)+t\ell_{t})$ variables. Note that most applications of the SDP to code-based cryptography, for instance in the McEliece scheme, choose $t\ll n$, hence the asymptotic bounds on the number of polynomials and variables in the improved modeling are both $\mathcal{O}(n\ell_{t})$. As shown in Table 1, our modeling improves over MPS by a factor of $\log_{2}(n)\log_{t}(n)$.
@@
Table 1: Comparison with the asymptotic size of the polynomial system in [34, Theorem 13], where $n$ is the length of the code and $t$ the bound on the weight of the target vector, that is $\mathsf{wt}(\mathbf{e})\leq t$.
@@ 5 Complexity Analysis of Modeling 2
We point out that the study in [13] investigates a particular case of the problem that this paper deals with, that is the regular syndrome decoding problem. The regular syndrome decoding problem considers error vectors having a regular distribution of non-zero entries. The algebraic modeling proposed in [13] is conjectured to exhibit semi-regular behavior when the linear parity-check constraints and the fixed, structured quadratic polynomials are considered separately. This suggests that, to some extent, their model behaves like a random polynomial system. Despite the fact that the problem tackled in [13] is a particular case of the problem we consider, our modeling has not been devised as a generalization of their modeling. Furthermore, we show that for the more general case, our modeling yields different results.
For the rest of this section, we retain the notation defined in Notation 2. We consider the polynomial ring $\mathbb{F}_{2}[X,Y]$ with the $\mathsf{degrevlex}$ term order with the $X$ variables greater than the $Y$ variables. Let $S\subset\mathbb{F}_{2}[X,Y]$ be the set of polynomials of Modeling 2 as described in Section 4.2. Let $L$ and $Q$ denote the sets of linear and quadratic polynomials, respectively. Clearly $S=L\cup Q$. Write also $L=L_{\mathbf{H}}\cup P$, where $L_{\mathbf{H}}$ denotes the set of linear polynomials in (2) introduced with the parity check matrix $\mathbf{H}$, and $P$ denotes the remaining linear polynomials in $S$. In other words, $P$ is the following set
$\begin{split}P=&\left\{x_{1}+y_{1,1},y_{1,2},\ldots,y_{1,\ell_{t}}\right\}\cup\left\{x_{i}+y_{i,1}+y_{i-1,1}\mid i=2,\ldots,n\right\}\\ \cup&\left\{y_{n,j}+v_{j}\mid j=1,\ldots,\ell_{t}\right\}.\end{split}$
We want to estimate the degree of regularity of $S$. Since we do not know $L_{\mathbf{H}}$ a priori, we consider the set $S\setminus L_{\mathbf{H}}=Q\cup P$ and compute its degree of regularity. Indeed, we found that analyzing the degree of regularity or solving degree of the system with the linear equations (2) of $L_{\mathbf{H}}$ included was too challenging and unpredictable, as it heavily depends on the specific instance of the parity check matrix $\mathbf{H}$. For this reason, we chose to establish mathematical results for the system without $L_{\mathbf{H}}$, with the aim of providing a clearer foundation. Notice that the degree of regularity of $S\setminus L_{\mathbf{H}}=Q\cup P$ gives an upper bound to the degree of regularity of the whole system $S$ (see Remark 8).
We break down the problem by first computing the degree of regularity of $Q$ and then that of $Q\cup P$. We take advantage of the fact that the Hilbert series of $Q$ and of $Q\cup P$ are polynomials and compute their degree, i.e. for instance, $d_{\mathrm{reg}}(Q)=\deg\mathrm{HS}_{\mathbb{F}_{2}[X,Y]/\langle Q^{\mathrm{top}}\rangle}(z)+1$ as per Remark 3, similarly for $Q\cup P$. To this end, we are going to compute the maximum degree of a monomial in $\mathbb{F}_{2}[X,Y]/\langle Q^{\mathrm{top}}\rangle$, similarly we do for $Q\cup P$.
@@
The ideal $\langle Q^{\mathrm{top}}\rangle\subseteq\mathbb{F}_{2}[X,Y]$ is thus a monomial ideal.
@@
  • If $n$ is even, there are exactly two sets of maximal cardinality $\mathcal{S}_{1},\mathcal{S}_{2}\subseteq\mathcal{N}$ such that no set in $\mathcal{P}$ is a subset of $\mathcal{S}$.
  • If $n$ is odd, there is exactly one set of maximal cardinality $\mathcal{S}\subseteq\mathcal{N}$ such that no set in $\mathcal{P}$ is a subset of $\mathcal{S}$.
@@
We aim to find the number of sets of maximal cardinality $\mathcal{S}\subseteq\mathcal{N}$ such that no pair from $\mathcal{P}$ (i.e., no two consecutive elements) appears in $\mathcal{S}$. In order to avoid pairs of consecutive elements, we can only select non-consecutive elements from $\mathcal{N}$. To maximize the size of $\mathcal{S}$, we select every other element from $\mathcal{N}$. The size of such a set of maximal cardinality $\mathcal{S}$ is: $\left\lceil\frac{n}{2}\right\rceil$. Thus:
  • If $n$ is even, a set of maximal cardinality contains $\frac{n}{2}$ elements.
  • If $n$ is odd, a set of maximal cardinality contains $\frac{n+1}{2}$ elements.
@@
    Start with 1 and select every other element: $\mathcal{S}_{1}=\{1,3,5,\dots,n-1\}.$ This set contains all the odd-numbered elements of $\mathcal{N}$, and its size is $k$.
@@
    Start with 2 and select every other element: $\mathcal{S}_{2}=\{2,4,6,\dots,n\}.$ This set contains all the even-numbered elements of $\mathcal{N}$, and its size is also $k$.
Since there are no other ways to select $k$ elements without picking consecutive elements, these are the only two sets of maximal cardinality for $n$ even.
Case 2: $n$ is odd. Let $n=2k+1$. The largest possible set $\mathcal{S}$ contains $k+1=\frac{n+1}{2}$ elements. In this case, there is only one way to construct a set of size $k+1$ that avoids consecutive elements, i.e. start with 1 and select every other element: $\mathcal{S}_{1}=\{1,3,5,\dots,n\}.$ This set contains $k+1$ elements and avoids consecutive pairs. If we were to start with 2 and select every other element, we would only get $k$ elements: $\mathcal{S}_{2}=\{2,4,6,\dots,n-1\}.$ This is not maximal, as it contains fewer than $k+1$ elements. Thus, for $n$ odd, there is exactly one maximal set. ∎
Lemma 2 can be used to prove the following corollary, which we will use to construct a maximal degree monomial in $\mathbb{F}_{2}[X,Y]/\langle Q^{\mathrm{top}}\rangle$. The idea behind the construction lies in the observation that a Gröbner basis of $Q^{\mathrm{top}}$ can be written as the union of disjoint subsets $Q^{\mathrm{top}}_{j,n}$ for $j=1,\ldots,\ell_{t}$, see Theorem 5.1, which we describe in the next corollary. Also, the next corollary computes a maximal degree monomial with respect to $Q^{\mathrm{top}}_{j,n}$ for every $j=1,\ldots,\ell_{t}$. Given these monomials, computing a maximal degree monomial in $\mathbb{F}_{2}[X,Y]/\langle Q^{\mathrm{top}}\cup P^{\mathrm{top}}\rangle$, or equivalently, the degree of its Hilbert series, becomes feasible with a slight modification of the subsets due to the presence of linear polynomials in $P^{\mathrm{top}}$.
@@
Let $n=8$ and $\ell_{t}=3$. According to Theorem 5.1 the degree of regularity of $Q$ is $21$. Using MAGMA, we compute and report the Hilbert series of the quotient ring $\mathbb{F}_{2}[X,Y]/\langle Q^{\mathrm{top}}\rangle$, i.e.
@@
$\tilde{Q}^{\mathrm{top}}_{j,n}:=Q^{\mathrm{top}}_{j,n-1}\setminus\{y_{1,j}y_{2,j}\}\subset\mathbb{F}_{2}[y_{2,j},\ldots,y_{n-1,j}],$
for every $j=1,\ldots,\ell_{t}$. Let $G$ be a Gröbner basis of $\langle Q^{\mathrm{top}}\cup P^{\mathrm{top}}\rangle$ as in Lemma 3. Due to the presence of the linear monomials contributed by $P^{\mathrm{top}}$, we observe
@@
Let $n=8$ and $\ell_{t}=3$. According to Theorem 5.2 the degree of regularity of $Q\cup P$ is $11$. Using MAGMA, we compute and report the Hilbert series of the quotient ring $\mathbb{F}_{2}[X,Y]/\langle Q^{\mathrm{top}}\cup P^{\mathrm{top}}\rangle$, i.e.